OpenStudy (anonymous):

w=f(z)=|z|^2-----------is it a continuity function or not?what is the derivative of this function? (z=complex number)---(Medal!)

OpenStudy (dan815):

@FutureMathProfessor

OpenStudy (zarkon):

The continuity part is easy...for the derivative use the Cauchy–Riemann equations

OpenStudy (anonymous):

Suppose z = x + i y then \[ f(z)=|z|^2= x^2 +y^2 + 0 i \] f takes only real values. Can be it be analytic?